OpenStudy (anonymous):

8 22/27 + 7 5/7 + 17 1/9

OpenStudy (anonymous):

?

OpenStudy (radar):

8 22/27 + 17 1/9 do these first: 8+17= 25 22/27 + 1/9=22/27 + 3/27=25/27 that sum for those 2 mixed fractions is 25 25/27now add the remainig term 7 5/7 25 + 7 =32 25/27 + 5/7 =175/189 +135/189 = 310/189= 1 121/189 32 +1 121/189= 33 121/189 33 121/189